# Astra[Uncaptioned image]: General Interactive World Model with Autoregressive Denoising

### v2

@@ 1 Introduction
Building generative world models is an emerging field where the ability to synthesize realistic and coherent video trajectories serves as a proxy for understanding and simulating the underlying dynamics of the world. With the rapid advances in visual generation (rombach2022ldm; blattmann2023stable-video-diffusion; yang2024cogvideox; brooks2024sora; wan2025wan), numerous video generation models have emerged and can perceive contextual cues and synthesize high-fidelity videos of open-world scenarios. These advances serve as the foundation for broader world simulation tasks, including game engines, autonomous driving, and spatial intelligence.
Standard text-to-video (T2V) or image-to-video (I2V) models typically produce only short, self-contained video clips conditioned on prompts or reference images. They lack the ability to generate coherent long-horizon rollouts that respond adaptively to external stimuli such as agent movements, viewpoint changes, or control signals. Without this responsiveness, these models fall short of simulating the interactive and causal dynamics of the real world. Furthermore, existing video generators are constrained by the finite temporal window of diffusion models, which prevents them from producing extended video sequences. Although recent works (mao2025yume; he2025matrix; song2025history; teng2025magi) have explored video continuation techniques or hybrid frameworks that combine autoregression with diffusion, they often struggle to strike a balance between maintaining consistency with historical frames and remaining responsive to new inputs. In addition, the autoregressive generation process introduces error accumulation, leading to degraded quality and coherence in long-term predictions. As a result, despite impressive progress in generative fidelity, existing approaches remain largely passive—capable of rendering visually compelling content yet lacking the interactivity, adaptability, and robustness required for true world simulation.
@@
We conduct extensive experiments on diverse open-source benchmarks, including Sekai (li2025sekai), SpatialVID (wang2025spatialvid), RT-1 (brohan2022rt), nuScenes (caesar2020nuscenes) and Multi-Cam Video (bai2025recammaster). As illustrated in Figure 1, Astra achieves state-of-the-art performance in action-driven video prediction, generating sequences that are highly interactive while maintaining visual coherence and dynamic consistency. Furthermore, our framework also demonstrates strong generalization across tasks and environments, underscoring its potential as a foundation for next-generation visual world models.
@@ 2 Related Work
Video generation models. In recent years, denoising diffusion models (ho2020ddpm; song2020ddim) have become the dominant paradigm in generative modeling, celebrated for their high fidelity and controllability. Following their success in text-to-image (T2I) synthesis (rombach2022ldm; dhariwal2021diffusion), early video generation methods (blattmann2023stable-video-diffusion; ho2022vdm) extended diffusion models to the temporal domain, typically by inflating image-based UNets (ronneberger2015unet) with additional temporal layers. More recently, Sora (brooks2024sora) and following studies (zheng1open-sora; yang2024cogvideox; ma2024latte) pioneer high-quality, high-resolution text-to-video (T2V) diffusion models. These models commonly use the diffusion transformer (DiT) architecture (peebles2023dit) to capture complex spatial and temporal coherence. More recent models, including (song2025history) and (zhang2025packinginputframecontext), further improve long-horizon consistency. Beyond pure diffusion, hybrid frameworks have been proposed to reconcile long-range prediction with high-quality synthesis. By combining autoregression for temporal modeling and diffusion for local realism, approaches such as StreamingT2V (henschel2025streamingt2v) and MAGI (teng2025magi) achieve extended video rollouts. However, these methods still struggle with error accumulation across long sequences and offer limited responsiveness to external actions, leaving a gap between video generation and true interactive world simulation.
@@
Visual world models. Beyond video generation, visual world models aim to capture the causal dynamics of the environment, enabling agents to simulate future trajectories and interact with the world. Unlike text-to-video or image-to-video models that generate short clips conditioned on static inputs, visual world models explicitly incorporate history and actions, making them essential for tasks such as planning, control, and embodied intelligence. Several recent works demonstrate this trend. (wu2024ivideogpt) extends autoregressive video transformers to integrate actions and rewards, allowing agents to predict how future observations evolve in interactive environments. (wang2024worlddreamer) formulates world modeling as masked-token prediction in discrete latent space, supporting multimodal conditioning and open-ended environment simulation. (huang2025vid2world; he2025matrix) adapts pretrained video diffusion models into an autoregressive framework with causal action guidance, making them capable of controllable video prediction. In navigation contexts, (bar2025navigation) employs conditional diffusion-transformers to generate plausible future agent observations, facilitating planning in unfamiliar scenes. Meanwhile, (cen2025worldvla) introduces a joint vision-language-action (VLA) framework that learns to model both world states and agent behaviors in an autoregressive manner. At larger scales, frameworks such as (bruce2024genie; agarwal2025cosmos; liu2024LWM) demonstrate that scaling up transformer architectures and extending context windows significantly improve rollout fidelity and generalization across domains. Recently, (mao2025yume) uses a Masked Video Diffusion Transformer (MVDT) that masks input features to improve video quality. Astra instead adopts a noise-as-mask strategy that selectively degrades past visual context, prompting the model to better integrate action signals and follow user commands. While such approaches demonstrate the promise of world models for interactive environments, they still face error accumulation, temporal drift in long rollouts, and limited responsiveness to diverse actions. Overcoming these issues remains key to building reliable and scalable visual world models.
@@ 3.2 Interactive world modeling via autoregressive denoising
Modern video generation models (brooks2024sora; wan2025wan) have demonstrated remarkable progress in simulating realistic visual dynamics. These models benefit from large-scale pretraining, which enables them to implicitly acquire partial knowledge of 3D spatial perception, temporal dependencies, and even simple physical patterns such as motion and force. However, despite their impressive fidelity, these models still fall short in constructing real-world scenes that can be preserved, interacted and explored. This raises a key question: Are T2V models really world models? In our view, the defining property of a world model is interactivity—the ability to adapt generation dynamically in response to arbitrary action inputs at arbitrary moments. While diffusion-based models can be conditioned on global prompts or scene attributes, such conditioning mechanisms do not enable fine-grained, online interaction. To address this limitation, we turn to the autoregressive framework, which naturally supports stepwise prediction conditioned on both past observations and current actions. Unlike diffusion models that generate video in a single pass, autoregression allows for instantaneous responses to action inputs, enabling controllable and adaptive rollouts. By integrating this property with the generative power of denoising models, we design an autoregressive denoising framework that achieves both high-quality synthesis and interactive controllability. Although prior works have explored combining autoregression and denoising, reconstructing this hybrid paradigm for world modeling remains non-trivial. Beyond simply chaining autoregression and denoising, we must carefully define the observation–action interface and design mechanisms to balance the trade-off between long-term consistency and immediate responsiveness
@@ 3.3 History condition with noisy memory
After addressing the challenge of action control, we turn to another open problem: balancing the long-term temporal consistency with responsiveness to actions. Prior works have shown that generating coherent long videos requires conditioning on an extended history. However, we observe a trade-off: increasing the length of history improves temporal consistency but weakens action responsiveness. We refer to this phenomenon as visual inertia—the tendency of the model to rely heavily on past visual information while overlooking user actions. This arises because real-world datasets contain predominantly smooth motions, leading the model to prioritize continuity over abrupt, action-driven changes. To mitigate this inherent contradiction, we avoid naively shortening the conditioning horizon and instead seek a more elegant solution. Considering the asymmetry between dense visual inputs and sparse action signals, we propose to reduce the dominance of visual conditioning by introducing controlled corruption. Unlike (mao2025yume), which randomly masks visual tokens, we adopt a noise-as-mask strategy: injecting random noise into the conditioning video to degrade and blur its information content (Figure 3). This design offers two advantages. First, it requires no architectural modifications or additional learnable parameters in the denoising model. Second, by corrupting the visual context, it prevents the model from directly copying clean frames and forces it to integrate action cues into generation. The corruption noise is independent of diffusion noise, so inference can use clean historical frames. Through this training strategy, the model learns to balance reliance on action and history, thereby overcoming visual inertia. To further extend the effective history horizon, we adopt the compression approach of (zhang2025packinginputframecontext), which retains the first frame while compressing the intermediate history into compact visual tokens, preserving long-range temporal information without overwhelming the action signal.
@@ 4.1 Experiment setups
Datasets. To train our model, we leverage a diverse set of datasets covering autonomous driving, egocentric exploration, multi-camera rendering, and robot control, as shown in Table 1. Specifically, we use nuScenes (caesar2020nuscenes) for vehicle pose prediction, Sekai (li2025sekai) and SpatialVID (wang2025spatialvid) for large-scale in-the-wild videos with rich camera annotations, Multi-Cam Video (bai2025recammaster) for synthetic multi-view sequences, and RT-1 (brohan2022rt) (via Open X-Embodiment (o2024open)) for robot action trajectories. All videos are resized and cropped to 480p, and action annotations are temporally aligned by interpolating to every 4 frames, matching the temporal compression of the video VAE. Together, these datasets provide complementary action signals (vehicle, camera, and robot pose) that support unified action-aware world modeling. For evaluation, we construct Astra-Bench, a benchmark comprising 20 held-out samples from each dataset, designed to cover a diverse range of real-world scenarios.
@@
Dataset Action Scenario Size nuScenes (caesar2020nuscenes) Camera (7) Autonomous driving 850 Sekai (li2025sekai) Camera (12) Walking & drone view 50K SpatialVid (wang2025spatialvid) Camera (7) & keyboard / mouse In-the-wild videos 200K RT-1 (brohan2022rt) Robotic pose (7) Robot manipulation 9978 Multi-Cam Video (bai2025recammaster) Camera (12) Human motion 136K Total - - $\sim$ 397K (360 hours)
@@
Method Instruction Following $\uparrow$ Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Wan-2.1 (wan2025wan) 0.061 0.854 0.903 0.958 0.489 0.691 MatrixGame (he2025matrix) 0.268 0.916 0.928 0.981 0.441 0.748 Yume (mao2025yume) 0.652 0.936 0.938 0.985 0.523 0.741 Astra (Ours) 0.669 0.939 0.945 0.989 0.531 0.747
Training details. We initialize our model from the pre-trained video diffusion model (wan2025wan) and train it on 8 GPUs (80 GB each) with a per-GPU batch size of 1. Optimization is performed with AdamW (loshchilov2017decoupled) using a learning rate of 1e51e-5 for 30 epochs, requiring about 24 hours to converge. The training is conducted in the latent space of a 3D VAE. In pixel space, the number of condition frames is randomly sampled from [1, 128], while the number of target frames is fixed to 33. Additional implementation details are provided in Appendix A.
Metrics. Astra-Bench evaluates two key aspects of world models: visual quality and instruction following (camera motion tracking), using six fine-grained metrics. For instruction following, we assess whether generated videos accurately reflect intended walking directions and camera movements. While pose estimation tools such as MegaSaM (li2025megasam) can automate this process, inaccuracies in camera motion prediction and quantization errors limit their reliability. We therefore adopt human evaluation to ensure accurate assessment. We adopt metrics from VBench (huang2024vbench) for the remaining dimensions (i.e., subject consistency, background consistency, motion smoothness, aesthetic quality, and image fidelity). All test videos are generated at 480×832 resolution, 20 FPS, and 96 frames, using 50 inference steps for each model.
@@ 4.2 Main results
We evaluate our method on Astra-Bench and compare it against recent state-of-the-art video generation and world modeling approaches, including Wan-2.1 (wan2025wan), Matrix-Game (he2025matrix) and YUME (mao2025yume). As shown in Table 2, our method consistently outperforms baselines across all metrics, demonstrating strong advantages in both visual quality and action-conditioned responsiveness. Astra achieves consistently superior results, surpassing existing video generation and world modeling approaches in both fidelity and controllability. Our model produces videos with sharper details, smoother motion, and stronger temporal coherence (Figure 5), leading to higher scores on visual quality metrics such as subject and background consistency, motion smoothness, and overall aesthetic appeal. More importantly, it demonstrates a clear advantage in instruction following: human evaluations confirm that the generated trajectories more faithfully follow intended camera movements and action directions compared to prior methods. Importantly, while competing approaches often suffer from error accumulation and drift during long rollouts, Astra maintains stability across extended horizons (Figure 6), making it particularly suitable for real-world interactive scenarios where both high fidelity and reliable action following are critical.
@@ 4.3 Analysis
Action integration via ACT-Adapter. Our lightweight ACT-Adapter provides an efficient way to inject action features into the pre-trained flow transformer. Freezing most parameters while tuning only the adapter and attention layers ensures maximum reuse of generative knowledge. Ablation results (Table 3, cross attn. adapter) demonstrate that ACT-Adapter achieves stronger action-conditioned performance than the cross-attention adapter used in he2025matrix, confirming its effectiveness as a simple yet powerful extension for enhancing interactivity.

### v3

@@ 1 Introduction
Building generative world models is an emerging field where the ability to synthesize realistic and coherent video trajectories serves as a proxy for understanding and simulating the underlying dynamics of the world. With the rapid advances in visual generation (Rombach et al., 2022; Blattmann et al., 2023; Yang et al., 2025; Brooks et al., 2024; Wan et al., 2025), numerous video generation models have emerged and can perceive contextual cues and synthesize high-fidelity videos of open-world scenarios. These advances serve as the foundation for broader world simulation tasks, including game engines, autonomous driving, and spatial intelligence.
Standard text-to-video (T2V) or image-to-video (I2V) models typically produce only short, self-contained video clips conditioned on prompts or reference images. They lack the ability to generate coherent long-horizon rollouts that respond adaptively to external stimuli such as agent movements, viewpoint changes, or control signals. Without this responsiveness, these models fall short of simulating the interactive and causal dynamics of the real world. Furthermore, existing video generators are constrained by the finite temporal window of diffusion models, which prevents them from producing extended video sequences. Although recent works (Mao et al., 2025; He et al., 2025; Song et al., 2025; Teng et al., 2025) have explored video continuation techniques or hybrid frameworks that combine autoregression with diffusion, they often struggle to strike a balance between maintaining consistency with historical frames and remaining responsive to new inputs. In addition, the autoregressive generation process introduces error accumulation, leading to degraded quality and coherence in long-term predictions. As a result, despite impressive progress in generative fidelity, existing approaches remain largely passive—capable of rendering visually compelling content yet lacking the interactivity, adaptability, and robustness required for true world simulation.
@@
We conduct extensive experiments on diverse open-source benchmarks, including Sekai (Li et al., 2025a), SpatialVID (Wang et al., 2025), RT-1 (Brohan et al., 2022), nuScenes (Caesar et al., 2020) and Multi-Cam Video (Bai et al., 2025). As illustrated in Figure 1, Astra achieves state-of-the-art performance in action-driven video prediction, generating sequences that are highly interactive while maintaining visual coherence and dynamic consistency. Furthermore, our framework also demonstrates strong generalization across tasks and environments, underscoring its potential as a foundation for next-generation visual world models.
@@ 2 Related Work
Video generation models. In recent years, denoising diffusion models (Ho et al., 2020; Song et al., 2020) have become the dominant paradigm in generative modeling, celebrated for their high fidelity and controllability. Following their success in text-to-image (T2I) synthesis (Rombach et al., 2022; Dhariwal and Nichol, 2021), early video generation methods (Blattmann et al., 2023; Ho et al., 2022) extended diffusion models to the temporal domain, typically by inflating image-based UNets (Ronneberger et al., 2015) with additional temporal layers. More recently, Sora (Brooks et al., 2024) and following studies (Zheng et al., 2024; Yang et al., 2025; Ma et al., 2024) pioneer high-quality, high-resolution text-to-video (T2V) diffusion models. These models commonly use the diffusion transformer (DiT) architecture (Peebles and Xie, 2023) to capture complex spatial and temporal coherence. More recent models, including (Song et al., 2025) and (Zhang and Agrawala, 2025), further improve long-horizon consistency. Beyond pure diffusion, hybrid frameworks have been proposed to reconcile long-range prediction with high-quality synthesis. By combining autoregression for temporal modeling and diffusion for local realism, approaches such as StreamingT2V (Henschel et al., 2025) and MAGI (Teng et al., 2025) achieve extended video rollouts. However, these methods still struggle with error accumulation across long sequences and offer limited responsiveness to external actions, leaving a gap between video generation and true interactive world simulation.
@@
Visual world models. Beyond video generation, visual world models aim to capture the causal dynamics of the environment, enabling agents to simulate future trajectories and interact with the world. Unlike text-to-video or image-to-video models that generate short clips conditioned on static inputs, visual world models explicitly incorporate history and actions, making them essential for tasks such as planning, control, and embodied intelligence. Several recent works demonstrate this trend. (Wu et al., 2024) extends autoregressive video transformers to integrate actions and rewards, allowing agents to predict how future observations evolve in interactive environments. (Wang et al., 2024) formulates world modeling as masked-token prediction in discrete latent space, supporting multimodal conditioning and open-ended environment simulation. (Huang et al., 2025; He et al., 2025) adapts pretrained video diffusion models into an autoregressive framework with causal action guidance, making them capable of controllable video prediction. In navigation contexts, (Bar et al., 2025) employs conditional diffusion-transformers to generate plausible future agent observations, facilitating planning in unfamiliar scenes. Meanwhile, (Cen et al., 2025) introduces a joint vision-language-action (VLA) framework that learns to model both world states and agent behaviors in an autoregressive manner. At larger scales, frameworks such as (Bruce et al., 2024; Agarwal et al., 2025; Liu et al., 2024) demonstrate that scaling up transformer architectures and extending context windows significantly improve rollout fidelity and generalization across domains. Recently, (Mao et al., 2025) uses a Masked Video Diffusion Transformer (MVDT) that masks input features to improve video quality. Astra instead adopts a noise-as-mask strategy that selectively degrades past visual context, prompting the model to better integrate action signals and follow user commands. While such approaches demonstrate the promise of world models for interactive environments, they still face error accumulation, temporal drift in long rollouts, and limited responsiveness to diverse actions. Overcoming these issues remains key to building reliable and scalable visual world models.
@@ 3.2 Interactive world modeling via autoregressive denoising
Modern video generation models (Brooks et al., 2024; Wan et al., 2025) have demonstrated remarkable progress in simulating realistic visual dynamics. These models benefit from large-scale pretraining, which enables them to implicitly acquire partial knowledge of 3D spatial perception, temporal dependencies, and even simple physical patterns such as motion and force. However, despite their impressive fidelity, these models still fall short in constructing real-world scenes that can be preserved, interacted and explored. This raises a key question: Are T2V models really world models? In our view, the defining property of a world model is interactivity—the ability to adapt generation dynamically in response to arbitrary action inputs at arbitrary moments. While diffusion-based models can be conditioned on global prompts or scene attributes, such conditioning mechanisms do not enable fine-grained, online interaction. To address this limitation, we turn to the autoregressive framework, which naturally supports stepwise prediction conditioned on both past observations and current actions. Unlike diffusion models that generate video in a single pass, autoregression allows for instantaneous responses to action inputs, enabling controllable and adaptive rollouts. By integrating this property with the generative power of denoising models, we design an autoregressive denoising framework that achieves both high-quality synthesis and interactive controllability. Although prior works have explored combining autoregression and denoising, reconstructing this hybrid paradigm for world modeling remains non-trivial. Beyond simply chaining autoregression and denoising, we must carefully define the observation–action interface and design mechanisms to balance the trade-off between long-term consistency and immediate responsiveness
@@ 3.3 History condition with noisy memory
After addressing the challenge of action control, we turn to another open problem: balancing the long-term temporal consistency with responsiveness to actions. Prior works have shown that generating coherent long videos requires conditioning on an extended history. However, we observe a trade-off: increasing the length of history improves temporal consistency but weakens action responsiveness. We refer to this phenomenon as visual inertia—the tendency of the model to rely heavily on past visual information while overlooking user actions. This arises because real-world datasets contain predominantly smooth motions, leading the model to prioritize continuity over abrupt, action-driven changes. To mitigate this inherent contradiction, we avoid naively shortening the conditioning horizon and instead seek a more elegant solution. Considering the asymmetry between dense visual inputs and sparse action signals, we propose to reduce the dominance of visual conditioning by introducing controlled corruption. Unlike (Mao et al., 2025), which randomly masks visual tokens, we adopt a noise-as-mask strategy: injecting random noise into the conditioning video to degrade and blur its information content (Figure 3). This design offers two advantages. First, it requires no architectural modifications or additional learnable parameters in the denoising model. Second, by corrupting the visual context, it prevents the model from directly copying clean frames and forces it to integrate action cues into generation. The corruption noise is independent of diffusion noise, so inference can use clean historical frames. Through this training strategy, the model learns to balance reliance on action and history, thereby overcoming visual inertia. To further extend the effective history horizon, we adopt the compression approach of (Zhang and Agrawala, 2025), which retains the first frame while compressing the intermediate history into compact visual tokens, preserving long-range temporal information without overwhelming the action signal.
@@ 4.1 Experiment setups
Datasets. To train our model, we leverage a diverse set of datasets covering autonomous driving, egocentric exploration, multi-camera rendering, and robot control, as shown in Table 1. Specifically, we use nuScenes (Caesar et al., 2020) for vehicle pose prediction, Sekai (Li et al., 2025a) and SpatialVID (Wang et al., 2025) for large-scale in-the-wild videos with rich camera annotations, Multi-Cam Video (Bai et al., 2025) for synthetic multi-view sequences, and RT-1 (Brohan et al., 2022) (via Open X-Embodiment (O’Neill et al., 2024)) for robot action trajectories. All videos are resized and cropped to 480p, and action annotations are temporally aligned by interpolating to every 4 frames, matching the temporal compression of the video VAE. Together, these datasets provide complementary action signals (vehicle, camera, and robot pose) that support unified action-aware world modeling. For evaluation, we construct Astra-Bench, a benchmark comprising 20 held-out samples from each dataset, designed to cover a diverse range of real-world scenarios.
@@
Dataset Action Scenario Size nuScenes (Caesar et al., 2020) Camera (7) Autonomous driving 850 Sekai (Li et al., 2025a) Camera (12) Walking & drone view 50K SpatialVid (Wang et al., 2025) Camera (7) & keyboard / mouse In-the-wild videos 200K RT-1 (Brohan et al., 2022) Robotic pose (7) Robot manipulation 9978 Multi-Cam Video (Bai et al., 2025) Camera (12) Human motion 136K Total - - $\sim$ 397K (360 hours)
@@
Method Instruction Following $\uparrow$ Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Wan-2.1 (Wan et al., 2025) 0.061 0.854 0.903 0.958 0.489 0.691 MatrixGame (He et al., 2025) 0.268 0.916 0.928 0.981 0.441 0.748 Yume (Mao et al., 2025) 0.652 0.936 0.938 0.985 0.523 0.741 Astra (Ours) 0.669 0.939 0.945 0.989 0.531 0.747
Training details. We initialize our model from the pre-trained video diffusion model (Wan et al., 2025) and train it on 8 GPUs (80 GB each) with a per-GPU batch size of 1. Optimization is performed with AdamW (Loshchilov and Hutter, 2017) using a learning rate of 1e51e-5 for 30 epochs, requiring about 24 hours to converge. The training is conducted in the latent space of a 3D VAE. In pixel space, the number of condition frames is randomly sampled from [1, 128], while the number of target frames is fixed to 33. Additional implementation details are provided in Appendix A.
Metrics. Astra-Bench evaluates two key aspects of world models: visual quality and instruction following (camera motion tracking), using six fine-grained metrics. For instruction following, we assess whether generated videos accurately reflect intended walking directions and camera movements. While pose estimation tools such as MegaSaM (Li et al., 2025b) can automate this process, inaccuracies in camera motion prediction and quantization errors limit their reliability. We therefore adopt human evaluation to ensure accurate assessment. We adopt metrics from VBench (Huang et al., 2024) for the remaining dimensions (i.e., subject consistency, background consistency, motion smoothness, aesthetic quality, and image fidelity). All test videos are generated at 480×832 resolution, 20 FPS, and 96 frames, using 50 inference steps for each model.
@@ 4.2 Main results
We evaluate our method on Astra-Bench and compare it against recent state-of-the-art video generation and world modeling approaches, including Wan-2.1 (Wan et al., 2025), Matrix-Game (He et al., 2025) and YUME (Mao et al., 2025). As shown in Table 2, our method consistently outperforms baselines across all metrics, demonstrating strong advantages in both visual quality and action-conditioned responsiveness. Astra achieves consistently superior results, surpassing existing video generation and world modeling approaches in both fidelity and controllability. Our model produces videos with sharper details, smoother motion, and stronger temporal coherence (Figure 5), leading to higher scores on visual quality metrics such as subject and background consistency, motion smoothness, and overall aesthetic appeal. More importantly, it demonstrates a clear advantage in instruction following: human evaluations confirm that the generated trajectories more faithfully follow intended camera movements and action directions compared to prior methods. Importantly, while competing approaches often suffer from error accumulation and drift during long rollouts, Astra maintains stability across extended horizons (Figure 6), making it particularly suitable for real-world interactive scenarios where both high fidelity and reliable action following are critical.
@@ 4.3 Analysis
Action integration via ACT-Adapter. Our lightweight ACT-Adapter provides an efficient way to inject action features into the pre-trained flow transformer. Freezing most parameters while tuning only the adapter and attention layers ensures maximum reuse of generative knowledge. Ablation results (Table 3, cross attn. adapter) demonstrate that ACT-Adapter achieves stronger action-conditioned performance than the cross-attention adapter used in He et al. (2025), confirming its effectiveness as a simple yet powerful extension for enhancing interactivity.